A young photographer took out a loan of $6472 to buy a camera. Payments were due to be made at the end of each month for 36 months at a rate of 8.7% compounded monthly. How much did he have to pay periodically?
Answer #1

PV = 6472
Rate per month = 8.7%/12
Number of Months = 36

Monthly payments = PV/(1-(1+r)-n)/r = 6472/(1-(1+8.7%/12)-36)/(8.7%/12) = 208.07

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well